# The ELM Neuron: an Efficient and Expressive

### v1

@@ 4.3 Evaluating on complex and very long temporal dependency tasks
To test the extent of the ELM neuron’s ability to extract complex long-range dependencies, we use three modern machine learning tasks. The Sequential CIFAR-10 task involves solving the 10-way CIFAR-10 Krizhevsky et al. (2009) classification problem given a flattened sequence of $1024$ pixel values, corresponding to $32\times 32$ sized images Tay et al. (2021). The Pathfinder and Pathfinder-X tasks Linsley et al. (2018); Tay et al. (2021) require determining whether two dots are connected by a path in a 2D image (see Figure 6, but the images are flattened out to be a sequence of pixels before being fed into the model. For the Pathfinder-X task, the length of this sequence is $16384$, corresponding to a $128\times 128$ image (see Figure 6), making it a notoriously difficult long-range modeling task.
Our results are summarized in Table 1, where we compare the ELM neuron against several strong baselines. The model most comparable to ours is an LSTM with explicit long timescale-derived gating bias initialization Tallec and Ollivier (2018) (Chrono-LSTM). We also incorporate two self-attention-based models, including the Transformer Vaswani et al. (2017) and the Longformer Bellec et al. (2018). Finally, we include Structured State Space Sequence model Gu et al. (S4) and a Transformer leveraging exponential moving averages Ma et al. (2023) (Mega), the latter being the current state-of-the-art.
When comparing between recurrent cells (ELM Neuron vs. Chrono-LSTM), we find that both models solve the tasks, albeit the ELM Neuron substantially outperforms Chrono-LSTM on the Pathfinder-X task. The self-attention-based models trail further behind, with both Transformer and Longformer failing to solve the Pathfinder-X task Tay et al. (2021). S4 and Mega perform better, but they are both deep architectures; Mega features 6 layers of 128 and 256 memory cells and self-attention, compared to the ELM neuron with 150 memory units and less than $100$K parameters.
Table 1: Long Sequence Modeling Datasets
Model Accuracy
Sequential
CIFAR-10
Pathfinder Pathfinder-X
ELM Neuron (ours) 54.0 74.72 77.3
Chrono-LSTM Tallec and Ollivier (2018) 46.61 70.79 51.85
Transformer Vaswani et al. (2017) 42.44 71.4 FAIL
Longformer Bellec et al. (2018) 42.22 69.71 FAIL
S4 Gu et al. 87.26 86.05 88.1
Mega Ma et al. (2023) 90.44 96.01 97.98
The ELM neuron can solve long-range sequence modeling tasks. Results on three complex temporal dependency tasks. Surprisingly, the simple ELM architecture is able to solve the notoriously challenging Pathfinder-X task with sample length $>16$K. It scores higher than the transformer baselines and is only outperformed by big 6-staged architectures tuned specifically for these tasks (S4, Mega).
[Uncaptioned image] Figure 6: Pathfinder-X Samples. The task is to classify whether two dots are connected by a dashed line. Each $128\times 128$ image is flattened to a sequence in raster-scan order, yielding sequences of total length $16384$.
@@
Overall, the results suggest that the ELM neuron is capable of solving tasks with very long temporal dependencies and doing so without employing multistage structures. We hypothesize that assembling the ELM neurons into layered network structures might give it enough processing capabilities to catch up with the deep models but leave this investigation to future work.
5 Conclusion
@@
In this study, we introduced a biologically inspired recurrent cell, the Expressive Leaky Memory (ELM) neuron, and demonstrated its capability to fit the full spike-level input/output mapping of a high-fidelity biophysical neuron model. Unlike previous works that achieved this fit with millions of parameters, our a variant of our model only requires eight thousand, thanks to the careful design of the architecture leveraging appropriate implicit biases. Furthermore, the ELM neuron can effectively model a cortical neuron without making rigid assumptions about the number of memory states, unlike existing phenomenological models. It also imposes fewer presumptions about the complexity of the integration mechanism and the exact values of the timescales associated with each memory state.
We further scrutinized the implications of this design on various tasks, including some notoriously challenging ones in the machine learning literature. The ELM neuron was found to be competitive, a notable feat considering its biological motivation, a characteristic often lacking in similar successful models. However, we acknowledge the ongoing debate regarding the degree of "biological plausibility" required in intelligent systems. Biological systems, including neurons, are subject to a variety of ’wetware’ constraints that may not need to be reproduced.
@@
Despite this caveat, the promising performance of the ELM neuron on machine learning tasks suggests that we are beginning to incorporate the inductive biases that drive the development of more intelligent systems. Future research focused on connecting ELM neurons into a network could provide even more insights into the necessary and dispensable elements for building smarter machines.
Acknowledgments and Disclosure of Funding
@@
This work was supported by a Sofja Kovalevskaja Award from the Alexander von Humboldt Foundation. We acknowledge the support from the BMBF through the Tübingen AI Center (FKZ: 01IS18039A and 01IS18039B). AL, GM, and BS are members of the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 39072764.
@@ Appendix A Implementation Details
All computations were performed using Python 3.9, and the following libraries were instrumental in our implementation: jax 0.3.14 (coupled with jaxlib 0.3.10 as a GPU back-end) for auto-grad and auto-vectorization; equinox 0.8.0, a jax-based neural network library; optax 0.1.3, a jax-based optimizer library; and pytorch 1.12.1 for data-loading. The code used for our experiments will be accessible at the public repository upon publication. To maintain anonymity, the essential excerpts of the code are presented in Section E.
Table S1: The default ELM neuron parameters and recommendations.
Hyper-Parameter NeuronIO Tuning Recommendation
$\boldsymbol{s_{0}}$ 0 /
$\boldsymbol{w_{s}}$ 0.5 /
$\boldsymbol{w_{s}}$ (bounds) > 0 >= 0
$\boldsymbol{\tau_{s}}$ 5ms greater for long and sparse data
$\boldsymbol{m_{0}}$ 0 /
$d_{m}$ $20$ between $10-250$
$\boldsymbol{\tau_{m}}$ (init technique) equally spaced evenly spaced on logarithmic scale
$\boldsymbol{\tau_{m}}$ (init range) 1ms,100ms 1ms to sample length in ms
$\boldsymbol{\tau_{m}}$ (bounds) 0ms,500ms same as init range or larger
@@
$d_{\mathrm{mlp}}$ $2*{\color[rgb]{0.00390625,0.44921875,0.69921875}d_{m}}$ $2*{\color[rgb]{0.00390625,0.44921875,0.69921875}d_{m}}$
$\lambda$ 10 between $1-10$
$\boldsymbol{w_{y}}$ (bias) True /
$\boldsymbol{w_{y}}$ (init technique) Kaiming Uniform /
$d_{\mathrm{tree}}$ 45 receive all input 2-5 times over
$d_{\mathrm{brch}}$ 100 $2*{\color[rgb]{0.3359375,0.70703125,0.9140625}d_{\mathrm{tree}}}$
@@
On a novel dataset, we recommend starting with a smaller $d_{m}$ around $100$, and small $\lambda$ around $3$, as increasing both simultaneously may lead to increased training instability. Potential future work could explore setting $\lambda$ automatically based on $d_{m}$, or learning it directly. The timescales $\boldsymbol{\tau_{m}}$ should generally be derived from the dataset length and the suspected timescales of the temporal dependencies within the data. Increasing $\boldsymbol{\tau_{s}}$ may help to enhance learning speed in some cases. When using the (branch) ELM it is important to sufficiently over-sample the input (more synapses than input) as significant expressivity stems from $\boldsymbol{w_{s}}$ doing the selection. Additional recommendations are summarized in Table S1.
@@ Appendix B Datasets and Training Details
For each task and dataset, the training dataset was deterministically split to create a consistent validation dataset, which was used for model selection during training and hyperparameter tuning. By default all models were trained using Backpropagation Through Time (BPTT) with a batch size of 8 and the Adam optimizer, employing a learning rate of $5e^{-4}$ and a cosine-decay schedule across the entire training duration. All experiments were run on a single A100-40GB or A100-80GB and ran less than 24h, Pathfinder-X being the notable exception.
@@
For training and evaluation the dataset was pre-processed in accordance with Beniaguev et al. [2021], by capping somatic membrane voltage at -55mV and subtracting a bias of -67.7mV. Additionally, the somatic membrane voltage was scaled by 1/10 for training. Training samples were 500ms long with a 1ms bin size and $\Delta t$. The ELM neuron used the default parameters from Table S1. Models were trained for 30 epochs of 11,400 batches per epoch using Binary Cross Entropy (BCE) for spike prediction and Mean Squared Error (MSE) for somatic voltage prediction, with equal weighting. Loss was calculated after a 150ms burn-in period. The mean and standard deviation over three runs is reported, with Root Means Squared Error (RMSE) and Area Under the Receiver Operator Curve (AUC) for voltage and spike prediction, respectively.
@@
The digits were preprocessed by cutting them to a uniform length of one second and binning the spikes using various bin sizes, the default being 10ms. The models were trained using the Adamax optimizer with a learning rate of $5e^{-3}$ for 70 epochs, for 814 and 2000 batches per epoch for SHD and SHD-Adding respectively, with $\Delta t$ set to the bin size, and dropout probability set to 0.5. The ELM used $\lambda=5$, $d_{m}=100$ and $\tau_{m}$ initialized evenly spaced between 1ms and 150ms with bounds of 0ms to 1000ms, and the LSTM used a hidden size of 250 and additional recurrent dropout of 0.3, while the SNN used a learning rate of $2e^{-3}$ no dropout but a $l1$ regularization on the spikes of $0.01$. Models were trained using the Cross-Entropy (CE) loss on the last float output of the respective model, and the performance was reported as prediction Accuracy, with mean and standard deviation calculated over five runs (chance performance being 1/19).
Long Sequence Modeling Datasets: The images were preprocessed by binning the 256 greyscale values into 16 values, and 32 different values for Pathfinder-X. The models were trained for 300 epochs with a batch size of 384, and batch size of 768 for Pathfinder-X. All trained using Cross-Entropy (CE) loss on the last output of the model, and the performance is reported as prediction Accuracy. The mean over three runs is reported for all experiments.
@@
Dataset Data Bins Batch Size LR $d_{m}$ $\lambda$ Timescales
CIFAR-10 16 384 $1e^{-3}$ 150 3 logspace: $1-10^{3}$
Pathfinder 16 384 $3e^{-4}$ 150 3 logspace: $1-10^{3}$
Pathfinder-X 32 768 $2e^{-4}$ 150 5 logspace: $1-10^{4}$
The ELM memory timescale bounds were matched to the initialization range. On the Pathfinder dataset an additional input dropout of $0.2$ was applied, however using $lr=2e^{-4}$ and no dropout still achieves $71.61\%$ accuracy. Furthermore, the ELM used a synapse tau of $150ms$ on the Pathfinder-X dataset, which we observed to increase the learning speed.
@@
Dataset Data Bins Batch Size LR Hidden Size Timescales
CIFAR-10 $16$ $384$ $1e^{-3}$ $150$ uniform: $1-2*10^{3}$
Pathfinder $16$ $384$ $1e^{-3}$ $150$ uniform: $1-2*10^{3}$
Pathfinder-X $32$ $768$ $1e^{-3}$ $70$ uniform: $1-2*10^{4}$
Alternative configurations for the Chrono-LSTM model, including learning rates of $2e^{-4}$, $5e^{-4}$, and $7e^{-4}$, as well as a hidden size of $150$ with half the batch size (reaching the 80GB GPU memory limit), were evaluated. For Pathfinder-X, only the reported configuration demonstrated performance exceeding chance levels, and only for one of the three trials.

### v3

@@ 4.3 Evaluating on complex and very long temporal dependency tasks
To test the extent and limits of the ELM neuron’s ability to extract complex long-range dependencies, we use the classic Long Range Arena (LRA) benchmark datasets Tay et al. (2021). It consists of classification tasks; three image-derived datasets Image, Pathfinder, and Pathfinder-X (images being converted to a grayscale pixel sequence), and three text-based datasets ListOps, Text, and Retrieval. Pixel and token sequences were encoded categorically, however, only considering 8 or 16 different grayscale levels for images. In particular, the Pathfinder-X task is notoriously difficult, as the task is to determine whether two dots are connected by a path in a $128\times 128$ image (~16k length).
Table 1: The ELM neuron can solve challenging long-range sequence modeling tasks. The table shows the mean accuracy on Long Range Arena (LRA) Benchmark Tay et al. (2021). The ELM neuron routinely scores higher than the Chrono-LSTM or the much larger Transformer or Longformer, and only the large multi-layered architectures tuned specifically for these tasks, such as S4 or Mega, outperform it. Surprisingly, it is also the only non purpose-built model that can reliably solve the notoriously challenging $16$K sample length Pathfinder-X task. Model sizes of the bottom baseline models are extracted from Gu et al. (2021)Ma et al. (2023)Tay et al. (2021). Training details and model hyper-parameters are detailed in Appendix Section B, and Tables S2 and S3.
Image Pathfinder Pathfinder-X ListOps Text Retrieval
ELM Neuron (ours) 49.62 71.15 77.29 46.77 80.3 84.93
Chr.-LSTM Tallec & Ollivier (2018) 46.09 70.79   $*{}^{*}$ 44.55 75.4 82.87
# parameters ~100k ~100k ~100k ~100k ~200k ~150k
Transformer Vaswani et al. (2017) 42.44 71.4 FAIL 36.37 64.27 57.46
Longformer Beltagy et al. (2020) 42.22 69.71 FAIL 35.63 62.85 56.89
S4 Gu et al. (2021) 87.26 86.05 88.1 58.35 76.02 87.09
Mega Ma et al. (2023) 90.44 96.01 97.98 63.14 90.43 91.25
# parameters ~600k ~600k ~600k ~600k ~600k ~600k
Our results are summarized in Table 1, where we compare the ELM neuron against several strong baselines. The model most comparable to ours is an LSTM with derived explicit gating bias initialization for effectively longer internal timescales Tallec & Ollivier (2018) (Chrono-LSTM). When comparing the two, we find that both models consistently perform well, except on the Pathfinder-X***Only once during hyper-parameter tuning did a single Chrono-LSTM run achieve barely above chance task which only the ELM can reliably solve, albeit using longer ${\color[rgb]{0.0078125,0.62109375,0.44921875}\bm{\tau_{s}}}$ than usual. The larger self-attention-based models trail further behind, with both Transformer Vaswani et al. (2017) and Longformer Beltagy et al. (2020) completely failing to solve the Pathfinder-X task Tay et al. (2021). Only the purpose-built architectures such as S4 Gu et al. (2021) and Mega Ma et al. (2023) (current SOTA) perform better, but they require many layers of processing and many more parameters than an ELM neuron, which uses 150 memory units and typically ~100k parameters.
@@
Overall, the results suggest that the simple ELM neuron architecture is capable of reliably solving challenging tasks with very long temporal dependencies. Crucially, this required using memory timescales initialized according to the task length and highly nonlinear synaptic integration into 150 memory units (See Appendix B). While the LRA benchmark revealed the single ELM neurons limits, we hypothesize that assembling ELM neurons into layered networks might give it enough processing capabilities to catch up with the deep models, but we leave this investigation to future work.
5 Discussion
@@
In this study, we introduced a biologically inspired recurrent cell, the Expressive Leaky Memory (ELM) neuron, and demonstrated its capability to fit the full spike-level input/output mapping of a high-fidelity biophysical neuron model (NeuronIO). Unlike previous works that achieved this fit with millions of parameters, a variant of our model only requires a few thousand, thanks to the careful design of the architecture exploiting appropriate inductive biases. Furthermore, unlike existing neuron models, the ELM can effectively model neuron without making rigid assumptions about the number of memory states and their timescales, or the degree of nonlinearity in its synaptic integration.
We further scrutinized the implications and limitations of this design on various long-range dependency datasets, such as a biologically-motivated neuromorphic dataset (SHD-Adding), and some notoriously challenging ones from the machine learning literature (LRA). Leveraging slowly decaying memory units and highly nonlinear dendritic integration into multiple memory units, the ELM neuron was found to be quite competitive, in particular, compared to classic RNN architectures like the LSTM, a notable feat considering its much simpler architecture and biological inspiration.
It should be noted that despite its biological motivation, our model cannot give mechanistic explanations of neural computations as biophysical models do, and that the task of fitting another neuron’s I/O is not itself a biologically relevant task for a neuron. Many biological implementation details are abstracted away in favor of computational efficiency and conceptual insight, and the required/recovered ELM neuron hyper-parameters depend on what constitutes a sufficiently good fit and the model’s subsequent use case. Furthermore, ELM is trained using BPTT, which is not considered biologically plausible in itself, and ELM learning likely relies on neuronal plasticity beyond synapses, the extent of which in biological neurons is still a subject of debate. Additionally, our neuron model dendrites are rudimentary (e.g., lacking apical vs basal distinction) and rely on oversampling synaptic inputs for performance so far. Finally, given the use of biologically implausible BPTT as a training technique and comparatively larger ELM neuron sizes on the later datasets, one should be careful to directly draw conclusions about the learning capabilities of individual biological cortical neurons.
@@
Despite these caveats, the ELM’s ability to efficiently fit cortical neuron I/O and its promising performance on machine learning tasks suggests that we are beginning to incorporate the inductive biases that drive the development of more intelligent systems. Future research focused on connecting smaller ELM neurons into larger networks could provide even more insights into the necessary and dispensable elements for building smarter machines.
Acknowledgments
@@
This work was supported by a Sofja Kovalevskaja Award from the Alexander von Humboldt Foundation. We acknowledge the support from the BMBF through the Tübingen AI Center (FKZ: 01IS18039A and 01IS18039B). AL, GM, and BS are members of the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 39072764. AS would like to thank the Max Planck Society for their generous financial support throughout the project. We would like to thank Antonio Orvieto for help with table S6. We thank the Max Planck Computing and Data Facility (MPCDF) staff, as the majority of computations were performed on the HPC system Raven.
@@ Appendix A Implementation Details
All computations were performed using Python 3.9, and the following libraries were instrumental in our implementation: jax 0.3.14 (coupled with jaxlib 0.3.10 as a GPU back-end) for auto-grad and auto-vectorization; equinox 0.8.0, a jax-based neural network library; optax 0.1.3, a jax-based optimizer library; and pytorch 1.12.1 for data-loading. The accompanying git repository for the project can be found under: https://github.com/AaronSpieler/elmneuron.
Table S1: The ELM neuron parameters and recommendations.
Hyper-Parameter NeuronIO Recommendation
$\bm{s_{0}}$ 0 /
$\bm{w_{s}}$ 0.5 /
$\bm{w_{s}}$ (bounds) > 0 >= 0
$\bm{\tau_{s}}$ all 5ms $\propto$ time sparse data (tune)
$\bm{m_{0}}$ 0 /
$d_{m}$ $20$ up to $250$ (primary TUNE)
$\bm{\tau_{m}}$ (init technique) equally spaced evenly spaced on log scale
$\bm{\tau_{m}}$ (init range) 1ms, 100ms 1ms, data length
$\bm{\tau_{m}}$ (bounds) 0ms, 500ms same as init range
@@
$d_{\mathrm{mlp}}$ $2*{\color[rgb]{0.00390625,0.44921875,0.69921875}d_{m}}$ /
$\lambda$ 10 5
$\bm{w_{y}}$ (bias) True /
$\bm{w_{y}}$ (init technique) Kaiming Uniform /
$d_{\mathrm{tree}}$ 45 $\propto$ inp. 2-5x sampled (tune)
$d_{\mathrm{brch}}$ 100 $\propto$ space sparse data (tune)
For all experiments $\bm{w_{y}}$, $w_{p}$ and $\bm{\tau_{m}}$ were learnable, with $\bm{w_{s}}$ crucially also learnable for Branch-ELM.
@@
We primarily recommend ablating $d_{m}$. In case of small $d_{m}$, exploring larger relative $d_{\mathrm{mlp}}$ might yield improved performance. For ELM with many small ${\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{\tau_{m}}}$ or larger ${\color[rgb]{0.87109375,0.5625,0.01953125}\lambda}$ we have observed spome trainig instability; seemingly resolved through modified memory update (see below). The timescales $\bm{\tau_{m}}$ should generally be derived from the dataset length and the suspected timescales of the temporal dependencies within the data; if reasonably initialized, learnability doesn’t seem to be necessary. Increasing $\bm{\tau_{s}}$ may help to enhance learning speed in case of temporally very sparse data. When using the Branch-ELM it is important to sufficiently over-sample the input (more synapses than inputs) as we suspect significant expressivity stemming from $\bm{w_{s}}$ doing the selection. Additional recommendations are summarized in Table S1.
Timescale parametrization:
The memory timescales ${\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{\tau_{m}}}$ are directly learnable model parameters. They are constrained to an apriori-specified bound, which is enforced through a $sigmoid$ rectification (the lower bound being $>0$). By defining ${\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{\kappa_{m}}}=\exp(\nicefrac{{-\Delta t}}{{{\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{\tau_{m}}}}})$, the resulting values are ensured to be within $[0,1]$ for all valid ${\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{\tau_{m}}}$, irrespective of $\Delta t$. In preliminary experiments we observed increased training stability as opposed to directly learning ${\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{\kappa_{m}}}$.
Improved implementation:
We found that enabling greater changes in ${\color[rgb]{0.87109375,0.5625,0.01953125}\bm{\Delta m_{t}}}$ by introducing a multiplicative factor ${\color[rgb]{0.87109375,0.5625,0.01953125}\lambda}$ improved training performance. Interestingly, the ${\color[rgb]{0.87109375,0.5625,0.01953125}\lambda}\cdot(1-{\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{\kappa_{m}}})$ term can then be seen as effectively using a $\lambda$ times faster input timescale than ${\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{\tau_{m}}}$. This notion can be explicitly implemented by substituting ${\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{\kappa_{\lambda}}}=\exp(-\Delta t/{\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{\tau_{\lambda}}})$ where ${\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{\tau_{\lambda}}}=\nicefrac{{{\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{\tau_{m}}}}}{{{\color[rgb]{0.87109375,0.5625,0.01953125}\lambda}}}$. The appoximation holds for ${\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{\tau_{m}}}>>{\color[rgb]{0.87109375,0.5625,0.01953125}\lambda}>1$, and only diverges for small ${\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{\tau_{m}}}$, where it allows for less stark ${\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{m}}$ changes than the original implementation. In preliminary experiments we found this modification to result in increased training stability for ELM neuron with many small ${\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{\tau_{m}}}$ or larger ${\color[rgb]{0.87109375,0.5625,0.01953125}\lambda}$ (see Figure S4).
Stable memory dynamics:
A necessary prerequisite for solving long-range credit assignment problems in recurrent architectures is to address the vanishing and exploding gradient problem. In the ELM neuron we achieve this by enforcing a stable dynamic of the memory units, which are primarily responsible for carrying information forward through time. The combination of controlled (slow) decay using ${\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{\kappa_{m}}}$ (addressing the vanishing gradient problem) and controlled (bounded) growth using the complementary $1-{\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{\kappa_{m}}}$ (addressing the exploding gradients problem), couples input to forget timescales using ${\color[rgb]{0.87109375,0.5625,0.01953125}\lambda}$ in a principled way such that ${\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{m}}$ will be bounded if $\Delta{\color[rgb]{0.87109375,0.5625,0.01953125}\bm{m}}$ is bounded (e.g. ensured through $tanh$ rectification), even if latter was generated using a highly nonlinear MLP (see Figure S9d). Note that in principle, this construction could also be applied to layer wise processing in depth, instead of in recurrent processing in time; however, we leave this experiment to future investigations.
The SNN implementation:
The LIF neuron based SNN consisted of an output layer ($N_{o}$ = "number of classes") and a recurrent layer ($N_{r}=500-N_{o}$), with $20\%$ of neurons being inhibitory. The spiking threshold was $v_{thr}=1$, and neurons were partially reset after firing using $v_{t+{\Delta t}}=v_{t}-0.9\cdot v_{thr}$. The membrane timescale was initialized to 25ms, and directly learnable like in the ELM neuron. Each of the 100 synaptic weights were initialized to $w_{s}=0.3/sqrt(100)$, and were rectified using $ReLU$. All neurons were randomly connected on a synapse-by-synapse basis with $90\%$ probability to the previous-layer, and $10\%$ probability to the own-layer. The output neurons output was low-pass filtered using constant 20ms, before being used by the cross-entropy function.
@@ Appendix B Datasets and Training Details
For each task and dataset, the training dataset was deterministically split to create a consistent validation dataset, which was used for model selection during training and hyperparameter tuning. All models were trained using Backpropagation Through Time (BPTT), and used a cosine-decay learning-rate schedule across the entire training duration of the training run. All experiments were run on a single A100-40GB or A100-80GB and ran less than 24h, Pathfinder-X being the notable exception.
@@
For training and evaluation the dataset was pre-processed in accordance with Beniaguev et al. (2021), by capping somatic membrane voltage at -55mV and subtracting a bias of -67.7mV. Additionally, the somatic membrane voltage was scaled by 1/10 for training. Training samples were 500ms long with a 1ms bin size and $\Delta t$. The ELM neuron used the default parameters from Table S1. Models were trained using the Adam optimizer with an initial learning rate of $5e^{-4}$ and a batch size of 8 for 30 epochs with 11,400 batches per epoch using Binary Cross Entropy (BCE) for spike prediction and Mean Squared Error (MSE) for somatic voltage prediction, with equal weighting. Loss was calculated after a 150ms burn-in period. The mean and standard deviation over three runs is reported, with Root Means Squared Error (RMSE) and Area Under the Receiver Operator Curve (AUC) for voltage and spike prediction, respectively. The model hyper-parameters and training settings were chosen based on validation RMSE in preliminary ablations.
@@
The digits were preprocessed by cutting them to a uniform length of one second and binning the spikes using various bin sizes, the default being 2ms. The models were trained using the Adamax optimizer with an initial learning rate of $5e^{-3}$ and a batch size of 8 for 70 epochs, with 814 or 2000 batches per epoch for SHD and SHD-Adding respectively, with $\Delta t$ set to the bin size, and dropout probability set to 0.5. The ELM and Branch-ELM used $\lambda=5$, $d_{m}=100$ and $\tau_{m}$ initialized evenly spaced between 1ms and 150ms with bounds of 0ms to 1000ms, whereas the LSTM used a hidden size of 250 and additional recurrent dropout of 0.3, while the SNN used a learning rate of $2e^{-3}$ no dropout but a $l1$ regularization on the spikes of $0.01$. The Branch-ELM over-sampled the input with $d_{\mathrm{tree}}=100$, $d_{\mathrm{brch}}=15$ and used random synapse to branch assignment. Models were trained using the Cross-Entropy (CE) loss on the last float output of the respective model, and the performance was reported as prediction Accuracy, with mean and standard deviation calculated over five runs (chance performance being 1/19). The model hyper-parameters and training settings were chosen based on validation Accuracy in preliminary ablations.
Long Range Arena Benchmark: The images-based datasets were preprocessed by binning the individual grey-scale values (256 total) into 16, or 8 for Pathfinder-X, different levels. For the text based datasets a simple one-hot token encoding was used. For the Retrieval task with a two-tower setup, the latent-dimension was 75 for both models. All models were trained using the Adam optimizer, with the ELM neuron using an initial learning rate of $42e-$ 4, and the the LSTM models working best with $31e-$ 3. All were trained using Cross-Entropy (CE) loss on the last output of the model, and the performance is reported as prediction Accuracy. The mean over three runs is reported for all experiments.
@@
Dataset Input Dim Batch Size Epochs $\tau_{s}$ Timescales
Image 16 384 300 5 logspace: $1-10^{3}$
Pathfinder 16 384 300 5 logspace: $1-10^{3}$
Pathfinder-X 8 768 300 150 logspace: $1-2*10^{4}$
ListOps 25 384 150 5 logspace: $1-2*10^{3}$
Text 169 384 150 5 logspace: $1-4*10^{3}$
Retrieval 105 384 150 5 logspace: $1-4*10^{3}$
The ELM memory timescale bounds were matched to the initialization range. The ELM used a synapse tau of $150ms$ on the Pathfinder-X dataset, which we observed to increase the learning speed significantly, however, smaller synapse tau can also work (e.g. see Figure S8). Otherwise, hyper-parameters were harmonized as much as possible, to demonstrate the robustness of the hyper-parameter choice.
@@
Dataset Input Dim Batch Size Epochs Timescales
Image $16$ $384$ 300 uniform: $1-2*10^{3}$
Pathfinder $16$ $384$ 300 uniform: $1-2*10^{3}$
Pathfinder-X $8$ $768$ 300 uniform: $1-2*10^{4}$
ListOps $25$ $384$ 150 uniform: $1-4*10^{3}$
Text $169$ $384$ 150 uniform: $1-8*10^{3}$
Retrieval $105$ $384$ 150 uniform: $1-8*10^{3}$
The Chrono-LSTM hyper-parameter tuning primarily concerned the learning rates and hidden sizes, however a learning rate of $1e^{-3}$ (among the tested) and a hidden size of $150$ (max tested) consistently performed best, the exception being Pathfinder-X, where during tuning a single run using a smaller hidden size performed slightly above change.